### v1

# Continual Adversarial Defense

Abstract
In response to the rapidly evolving nature of adversarial attacks on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that can generalize to all types of attacks, including unseen ones, is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks used by many attackers. The defense system needs to upgrade itself by utilizing few-shot defense feedback and efficient memory. Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage. In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images. We leverage cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of 10 modern adversarial attacks and significant improvements over 10 baseline methods. In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against old attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.  [^1] Code at https://github.com/cc13qq/CAD
Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender. During the $i$-th stage, the defender $f_{i}$ leverages few-shot defense feedback provided by the security department and users of the target model to adapt to new attacks. The defender also needs to efficiently use memory to store the cache, retaining knowledge of past attacks.
1 Introduction
In recent years, deep neural networks have been widely used in numerous vision tasks [30], leading to remarkable breakthroughs. However, many neural networks are vulnerable to adversarial attacks, which aim to deceive a classifier by adding subtle perturbations to input images, thus altering prediction results. This vulnerability seriously jeopardizes the reliability of deep neural networks, particularly in security- and trust-sensitive domains.
Researchers have proposed various general defense methods against adversarial attacks. As a standard defense method, adversarial training [44] aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream [41] to remove potential adversarial perturbations or noise that could deceive the model. However, models of adversarial training often exhibit reduced classification accuracy on adversarial examples and may sacrifice classification capacity on clean images. On the other hand, purification techniques may appear more reliable, but the denoising procedure can inadvertently smooth the high-frequency texture of clean images, potentially causing issues for the target model.
From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones. The truth is, the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32]. As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [16]. The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [10].
To simulate a dynamic environment more realistically, we propose a defense scenario that considers practicality. As illustrated in Fig. 1, the attacker employs diverse attacks at different stages to target the model, while the defender adapts to emerging attacks by leveraging few-shot defense feedback. In this paper, we use adversarial examples with ground truths as the defense feedback, which is provided by the security department and users of the target model in real-world situations. In this ongoing “arms race,” the attacker operates within the gray-box setting [32], where they possess knowledge of the classifier’s architecture and have access to training data but are unaware of the defense mechanism in place. On the other hand, the defender has the same training data and access to the target model, but also possesses full knowledge of the initial attack.
Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting. As defenders, it’s essential to adapt to various new attacks at different stages while retaining knowledge of previous ones. (2) Few-shot adaptation. Increased feedback indicates more successful attacks on the target model. Therefore, we cannot afford too many successful attacks before upgrading the defense. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks results in accumulating defense feedback, potentially leading to memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this. (4) High accuracy in classifying both clean and adversarial images. A robust defense should not compromise the interests of those it protects. Hence, maintaining high classification accuracy on both clean and adversarial images is crucial.
In this paper, we propose the Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense feedback and efficient memory. In the first stage, we train an initial defense model using adversarial images from the initial attack. This model is specifically designed to classify images with adversarial noise, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes. We then fine-tune this expanded layer using few-shot defense feedback. To address the over-fitting issue that arises from the few-shot fine-tuning process, we reserve embedding space in the defense model for future attacks. This is achieved by generating and assigning virtual prototypes, which help compress the embedding of previous attacks. To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any feedback explicitly. Simultaneously, we utilize a small model to ensemble the defense model with the target model by estimating reliable logits for input images, ensuring maximum classification accuracy for both clean and adversarial images.
Based on our comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot feedback while maintaining high accuracy on clean images. Our main contributions can be summarized as follows:
$\bullet$ In order to simulate the rapidly evolving adversarial attacks against image classifiers, we introduce, for the first time, a dynamic scenario of various attacks that emerge in different stages. The defender is required to adapt to new attacks using few-shot defense feedback and efficient memory, while also retaining knowledge of previous attacks and upholding high performance on both clean and adversarial images.
$\bullet$ We propose the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial images.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory, and demonstrate significant improvements over 10 baseline methods.
2 Related Work
2.1 Adversarial Attacks
Adversarial attacks [6] attempt to fool a classifier into changing prediction results by attaching subtle perturbations to the input images while maintaining imperceptibility from human eyes. PGD [26] and BIM [14] generate adversarial examples through several iterations of perturbation and get significantly improved attack performance. DIM [40], TIM [4], and SIM [17] improve the transferability under the black-box setting and breach several defense techniques. Incorporating variance tuning, VIM [38] achieves a high attack rate against multiple advanced defense methods. In this paper, a wide spectrum of attacks is used to offense the target model at different stages under the gray-box setting.
2.2 Adversarial Defense
Adversarial training [26, 45] is the mainstream of adversarial defense that uses attacked samples to train a robust model. TRADES [44] characterize the trade-off between accuracy and robustness for classification problems via decomposing the robust error as the sum of the natural error and the boundary error. JEM [7] uses energy-based training of the joint distribution and improves the calibration and robustness of target models. Replacing the denoising diffusion probabilistic model with the elucidating diffusion model, DMAT [39] improves adversarial training and achieves new SOTA robust accuracy.
Another technical solution for protecting DNNs from adversarial attacks is adversarial purification [29, 9], which aims to purify the data stream before feeding it into the target model, functioning under the gray-box setting [32]. Based on an EBM trained with Denoising Score-Matching, Yoon et al. [41] propose Adaptive Denoising Purification (ADP) which conducts purification in a randomized purification scheme. Nie et al. [27] propose DiffPure that uses the forward and reverse processes of diffusion models to purify adversarial images.
Certainly, both adversarial training and purification methods have shown limitations in robustness, leading to reduced classification performance on clean images. Furthermore, the dynamic nature of the environment introduces a diverse array of unique attacks employed by multiple attackers. In this paper, we propose Continual Adversarial Defense (CAD) framework as a solution to effectively combat various attacks emerging stage by stage, while maintaining strong performance in classifying clean images.
2.3 Continual Learning
Continual learning (CL) [21, 28, 19, 18, 20, 25, 22, 24, 46, 23] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks. Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [28], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [48]. In recent years, some methods have aimed to solve the CL problem without depending on preserving data (called none-exemplar) [51], and some methods have attempted to learn new classes in a few-shot scenario [43] in which only a small number of new class data is gainable. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense feedback and efficient memory for adaptation.
3 Threat Model and Defense Principles
The defense system operates within a dynamic environment characterized by a multitude of unique attacks since the growing number of defenses naturally encourages the invention of new attack algorithms. Once frustrated by an existing defense, attackers undoubtedly resort to deploying various attacks in pursuit of breakthroughs. In response to this reality, we simulate the dynamic environment and put forth a practical threat model. In this threat model, attackers deploy different attacks in various stages to assault the target model, and the defender must adapt to evolving attacks by utilizing the defense feedback, just like an ”arms race”. The defense feedback is typically provided by security services and users of the target model upon the discovery of successful attacks during screening or cleanup efforts in the real world, and plays a critical role in informing and shaping the defender’s response to emerging threats. In this paper, we use adversarial examples with ground truths as the defense feedback. We formulate this attack vs defense scenario as follows.
Our scenario is established on the $N$-way $K$-shot classification paradigm. The attacker attacks under the gray-box setting [32] in which the attacker knows the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, the attacker knows the failure of previous attacks and turns to a new attack $A_{i}(\cdot)$. Meanwhile, the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning. And then, a weight estimator model $f_{p}$ is trained for the model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator to maintain high accuracy of both clean and adversarial images.
Taking reality into account, the defense mechanism should satisfy the following principles:
Principle 1
Continual adaptation to new attacks without catastrophic forgetting.
In the dynamic environment, it is not uncommon for certain attackers to employ the same attack across different stages. Therefore, the defense mechanism must be capable of adapting to a range of new attacks in each stage while also retaining knowledge of previous ones.
Principle 2
Few-shot adaptation.
The increase in defense feedback directly correlates with the frequency of successful attacks on the target model. As defenders, we must proactively adapt the defense based on few-shot feedback to prevent potential security disasters that could arise from delayed action following an abundance of feedback.
Principle 3
Memory-efficient adaptation.
Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to the memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense feedback for adaptation is impractical and should be avoided.
Principle 4
High accuracy in classifying both clean and adversarial images.
The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation. This decline could potentially have a negative impact on crucial real-world business operations. As a result, it is imperative for a defender to prioritize maintaining performance on both clean and adversarial images.
4 Continual Adversarial Defense Framework
As shown in  Fig. 2, Continual Adversarial Defense (CAD) consists of continual adaptation to defend the target model while preventing catastrophic forgetting ( Sec. 4.1), embedding reservation to alleviate overfitting brought by the few-shot feedback ( Sec. 4.2), prototype augmentation for memory efficiency ( Sec. 4.3), and model ensemble to ensure the classification performance on both clean and adversarial images(Section 4.4).
4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$. Inspired by continual-learning (CL) [28], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
Following [21], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ at the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas:
$y^{i}=y+N\times{i}.$ (1)
The initial defense model $f_{0}$ will be trained in Sec. 4.2. After this, we freeze the feature extractor $\varphi$ and fine-tune the part of the classifier expanded for the new attack $A_{i}$ using few-shot feedback $\mathcal{A}_{\mathrm{train}}^{i}$ in Sec. 4.3.
For evaluation, we are just focusing on which class the image is placed in rather than being threatened by which attack. Thus the prediction for an instance $\mathbf{x}$ is:
$y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\%N$ (2)
4.2 Embedding Reservation for Few-Shot Adaptation
A problem brought by the few-shot feedback in Principle 2 is over-fitting which is common in few-shot learning [31]. To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ‘virtual classes [43], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
Second, virtual instances are constructed by manifold mixup [36]:
$\mathbf{z}=h_{2}(\lambda h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\lambda)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as [48]. $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
Third, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
$\mathcal{L}_{\mathrm{v}}=$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{ce}(f_{0}(\mathbf{z}),\hat{y})+$
$+$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}}).$ (4)
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss [47], and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
$\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})=f_{0}(\mathbf{x}_{\mathrm{adv}}^{0})\otimes(\mathbf{1}-\mathrm{onehot}(y^{0}))$ (5)
where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector. The final loss for training the initial defense model is:
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\quad\mathcal{L}_{0}=\mathcal{L}_{\mathrm{ce}}+\gamma\mathcal{L}_{\mathrm{virtual}}$ (6)
In Sec. 4.2, it forces an instance to be away from the reserved virtual classes (item 1) and pushes the mixed instance towards a virtual class (item 2) and away from other classes (item 3). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved [48]. The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot feedback.
4.3 Prototype Augmentation for Memory Efficiency
In response to Principle 3, we use prototype augmentation to efficiently use memory.
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50]. Many CIL methods store a fraction of old data to jointly train the model with current data to address this issue [28]. However, preserving old feedback may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [51] to maintain the decision boundary of previous stages, without saving any feedback. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
$\mathbf{p}_{B}=U_{B}(P_{e}),\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (7)
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ the same as ProtoNet [31]:
$\mathbf{p}_{e}^{j}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (8)
where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of feedback for each class.
In each stage, the defense model is adapted using the following loss:
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{proto}}$ (9)
Algorithm 1 Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:  Generate initial adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}$ using $A_{0}$.
2:  Train weight estimator $f_{p}$ on $\mathcal{D}_{\mathrm{train}}$ using self-perturbation.
3:  Rewrite the ground-truth of $\mathcal{A}_{\mathrm{train}}^{i}$ using Eq. 1.
4:  Optimize defense model $f_{0}$ on $\mathcal{A}_{\mathrm{train}}^{0}$ using Eq. 5.
5:  Generate prototypes of $A_{0}$ using Eq. 8 and add them to prototype set $P_{e}$.
6:  for $i$ in $1,\cdots,T$ do
7:     Fine-tune the part of the classifier preserved for the new attack using Eq. 9.
8:     Generate prototypes of $A_{i}$ using Eq. 8 and add them to prototype set $P_{e}$.
9:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq. 10.
10:  end for
4.4 Model Ensemble for Clean and Adversarial Image Classification
To maintain high classification accuracy on both clean and adversarial images mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) [34] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean images. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models
We adopt the self-perturbation [37] to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. Details of the self-perturbation are given in the supplementary.
The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (10)
In this way, both the clean images and the adversarial images are correctly classified.
The overall algorithm of CAD framework is presented in Algorithm 1.
5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from two research streams. For evaluation and analysis, we use two metrics in this section: (1) classification accuracy against each attack after adaptation. (2) the average of the accuracy of all attacks that have occurred in and before in each stage, i.e. average incremental accuracy [28].
Datasets
CIFAR-10 [13] is a widely used dataset for adversarial attack and defense, containing 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 [33] is a subset of ImageNet [1] and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attack algorithms
We select 11 adversarial attack methods PGD [26], FGSM [6], BIM [14], RFGSM [35], MIM [3], DIM [40], NIM [17], SNIM [17], VNIM [38], and VMIM [38] to compose the attack pool. The perturbation magnitude of all the attacks is set to $\epsilon=8$ under $l_{\infty}$ following [44].
Implementation details
Both the target model and defense model use ResNet as the backbone network in our method. Specifically, we use WideResNet-28-10 [42] for CIFAR-10 following [7] and ResNet-50 [8] for ImageNet-100 following [15]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [5] is adopted as the classifier of the defense model. We utilize Torchattacks [12] for generating adversarial images. The amount of feedback per class is set to $K=10$. The number of stages is set to $T=9$ since there are 10 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [48]. Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
Baselines
We first compare to adversarial training methods TRADES [44], JEM [7], GAIRAT [45], FastAdv [11], RPF [2] and DMAT [39]. Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In premium-CAD, we train additional attack-specific defense models using abundant feedback (i.e., 1000 images per class ) in each stage, and ensemble them with the target model by the weight estimator. Details about premium-CAD are given in the supplementary.
Stage 0 1 2 3 4 5 6 7 8 9 -
Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
TRADES [44] 55.8 57.4 60.7 55.9 55.8 57.1 57.1 58.9 60.2 57.3 84.8
JEM [7] 55.8 46.5 60.7 52.3 52.4 51.1 51.4 57.0 62.0 48.2 91.7
PAT [15] 45.5 43.8 49.8 45.5 45.3 45.2 45.2 43.9 45.1 45.4 69.3
GAIRAT [45] 66.9 66.4 68.3 66.9 66.7 66.5 66.5 68.6 70.4 66.4 89.4
FastAdv [11] 37.1 36.8 41.9 37.0 37.7 37.7 54.6 36.4 39.6 37.6 75.2
RPF [2] 60.9 52.2 58.0 58.9 54.9 57.6 57.5 53.5 52.8 52.6 83.5
DMAT [39] 71.7 73.1 75.8 71.7 71.7 73.0 73.0 75.5 76.0 73.0 92.4
EBM [9] 75.2 73.0 70.0 76.6 76.5 72.4 73.9 74.0 72.9 73.8 86.8
ADP [41] 85.4 79.3 82.0 84.4 85.1 84.1 83.7 81.0 80.5 79.1 80.5
DiffPure [27] 72.5 69.8 71.8 70.1 72.2 73.4 73.5 72.7 70.2 69.4 89.3
CAD(ours) 95.8 93.4 83.8 94.7 95.0 93.7 93.7 95.2 95.4 93.7 95.8
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.6 94.6 94.1 94.6 94.3 95.1 95.1 95.5 95.7 96.0
Table 1: Classification accuracy (%) against various attacks on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by the efficient memory and few-shot defense feedback.
Stage 0 1 2 3 4 5 6 7 8 9 -
Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
TRADES [44] 46.5 49.6 48.8 46.2 48.0 47.7 46.7 48.2 45.5 49.5 76.2
GAIRAT [45] 59.8 58.4 59.4 59.2 58.8 58.9 58.9 59.7 60.2 59.8 74.1
RPF [2] 56.4 51.3 55.8 56.6 54.4 52.6 53.9 53.3 49.9 50.1 69.4
DMAT [39] 64.3 66.2 66.4 65.9 65.4 64.4 64.1 65.8 65.1 65.6 75.9
ADP [41] 77.2 71.8 73.7 71.8 75.6 75.9 76.2 76.0 76.9 72.3 74.1
DiffPure [27] 53.2 52.4 53.7 55.8 53.9 53.4 52.6 54.7 53.0 52.3 73.6
CAD(ours) 83.7 82.0 76.2 82.5 83.5 82.0 81.9 82.2 83.4 81.9 88.1
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 84.0 84.1 82.6 84.3 83.8 84.2 82.2 83.9 81.4 87.8
Table 2: Classification accuracy (%) on ImageNet-100.
5.1 Comparisons to Baselines
We compared CAD to other baseline methods in defending against 10 adversarial attacks. In our approach, the defense environment is dynamic, with attacks occurring at different stages. In contrast, the other baseline methods operate in a static defense environment that does not differentiate the order of attacks. We report the robust accuracy, which measures the performance on adversarial examples generated by attacks, and the standard accuracy, which measures the performance on clean data, separately for each method. We compare the defense performance of our method to other methods on CIFAR-10 and ImageNet-100. All the evaluation is conducted on the whole test set.
As depicted in  Tab. 1 and Tab. 2, despite the modest scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ in CIFAR-10 and $6\%$ in ImageNet-100 for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy for CIFAR-10 and $88.1\%$ for ImageNet-100, while the testing accuracy of the target model stands at $96.4\%$ for CIFAR-10 and $89.8\%$ for ImageNet-100. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images. We observe that the accuracy against FGSM is relatively lower than that for other attacks, largely due to the distinct noise pattern of the one-step FGSM compared to other attacks.
To assess CAD’s performance in mitigating catastrophic forgetting, we present the average incremental accuracy in Fig. 3. The curve initially experiences a slight decline in the first three stages before stabilizing. This temporary dip is primarily attributable to FGSM’s influence, which pulls down the entire curve, and the adaptation to new attacks, resulting in a minor decrease in accuracy against previous attacks. Subsequently, the curve exhibits a modest increase due to the resemblance of other attacks to the initial attack PGD, thereby elevating the overall curve. Throughout the evaluation, the curves consistently hover around $92\%$ for CIFAR-10 and $80\%$ for ImageNet-10, underscoring CAD’s robust capacity to counter catastrophic forgetting.
Note that premium-CAD outperforms all baseline methods. Given adequate defense feedback and memory resources, premium-CAD emerges as the superior choice.
5.2 Ablation Study
Figure 3: Average incremental accuracy (%) of CAD.
As argued, embedding reservation allows the defense model to be adapted more easily in the future and alleviates overfitting, prototype augmentation maintains the decision boundary of previous stages, and model ensemble ensures robust classification performance for both clean and adversarial images. Subsequently, we proceed to undertake an ablation study to validate the efficacy of each component.
As shown in Tab. 3 and Fig. 4 (Left), CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.2\%$ and $2.2\%$ respectivily in performance for adaptation. Furthermore, the absence of model ensemble results in the defense model’s inability to uphold classification accuracy for clean images alone. Additionally, we present curves illustrating the classification accuracy of each attack in Fig. 4 (Right). These curves demonstrate that the adaptation of CAD to FGSM is comparatively less effective than for other attacks, aligning with our earlier observations about FGSM in Sec. 5.1.
Stage 0 1 2 3 -
Attack PGD VNIM FGSM BIM Clean
w/o $L_{\mathrm{virtual}}$ 95.6 86.6 70.6 78.5 95.8
w/o $L_{\mathrm{proto}}$ 95.7 93.7 81.6 88.2 95.8
w/o $f_{p}$ 95.8 93.4 83.8 94.7 55.9
CAD(ours) 95.8 93.4 83.8 94.7 95.8
Table 3: Classification accuracy (%) on CIFAR-10 for ablation.
Figure 4: Accuracy (%) in each stage on CIFAR-10. Left: average incremental accuracy of ablation. Right: classification accuracy against each attack.
5.3 Multiple Attacks in One Stage
In practice, it is common for multiple attacks to occur simultaneously within a single stage. To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks. As presented in Tab. 4, when the number of attack increase to $2$ and $3$, the classification accuracy against these attacks still remain at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results presented in Tab. 4 indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
Stage 0 1 2 3
$\mathrm{number}=1$ 95.8 93.4 83.8 94.7
$\mathrm{number}=2$ 95.8 88.5 93.5 92.8
$\mathrm{number}=3$ 95.8 89.5 93.6 92.5
Table 4: Classification accuracy (%) on CIFAR-10 for different number of attacks in one stage.
5.4 Number of Shots for Adaptation
As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we conducted tests with $K=15,10,5,1$. The results presented in Tab. 5 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($90.2\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
Stage 0 1 2 3
Attack PGD VNIM FGSM BIM
$K=15$ 95.8 93.7 81.2 94.8
$K=10$ 95.8 93.4 83.8 94.6
$K=5$ 95.8 93.2 80.8 93.7
$K=1$ 95.8 92.8 77.5 90.2
Table 5: Classification accuracy (%) on CIFAR-10 with different feedback scale $K$.
5.5 Memory Allowance for Adaptation
As outlined in Principle 3, we need to employ memory efficiently for continual defense. In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each stage. As depicted in Tab. 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
Cache Size Model Size
Stage 1 2 3 1 2 3
CAD 50.3K 75.5K 100.4K 141M 141M 141M
$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD on CIFAR-10.
6 Conclusion
The operational environment for defense systems is inherently dynamic, and it is unrealistic to expect a defense method to address all types of attacks. As a response to this situation, we propose the first continual adversarial defense (CAD) framework, designed to dynamically adapt to a wide array of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles. Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
7 Acknowledgements
This work was supported in part by the Natural Science Foundation of China under Grant 61972169,62372203 and 62302186, in part by the National key research and development program of China(2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042), in part by the Knowledge Innovation Program of Wuhan-Basic Research, in part by China Postdoctoral Science Foundation 2022M711251.
Supplementary Material
A Weight Estimator Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
Algorithm 2 Self-perturbation
Input: A empty perturbation matrix $\boldsymbol{\eta}^{p}\in\mathbb{R}^{H\times W\times 3}$ with the same shape of real image $\mathbf{x}^{r}$.
Parameter: Max perturbation magnitude $\epsilon$, pattern mode.
Output: Self-perturbed image $\mathbf{x}^{p}$.
1:   A random direction matrix $R=\{\vec{r}_{ij}\}\in\mathbb{R}^{H\times W\times 3}$.
2:   for $\vec{\eta}_{ij}$ in $\boldsymbol{\eta}^{p}$ do
3:      Select random noise value $\alpha$.
4:      if pattern mode is ‘point-wise’ then
5:         $\vec{\eta}_{ij}:=\alpha\cdot\vec{r}_{ij}$.
6:      else if pattern mode is ‘block-wise’ then
7:         Select a random neighborhood $A_{ij}$ of $\vec{\eta}_{ij}$.
8:         for $\vec{\eta}_{ijk}$ in $A_{ij}$ do
9:            $\vec{\eta}_{ijk}:=\alpha\cdot\vec{r}_{ij}$.
10:         end for
11:      end if
12:   end for
13:   Clip perturbation $\boldsymbol{\eta}^{p}$ using Equation 11.
14:   Generate self-perturbed image $\mathbf{x}^{p}$ using Equation 12.
15:   return $\mathbf{x}^{p}$
Self-Perturbation
As presented in Algorithm 2, we perturb each point in the point-wise pattern and each block in the block-wise pattern in a stochastic direction, where blocks are random neighborhoods of a set of scattered points. The generated perturbation image $\boldsymbol{\eta}^{p}$ is constrained in $l^{\infty}$ norm, and clipped according to $\epsilon$,
$\boldsymbol{\eta}^{p}=\mathrm{Clip}_{[-\epsilon,\epsilon]}(\boldsymbol{\eta}^{p}).$ (11)
A self-perturbed image is calculated as
$\mathbf{x}^{p}=\mathbf{x}^{r}+\boldsymbol{\eta}^{p}.$ (12)
Model Ensemble
Trained with self-perturbation, the weight estimator $f_{p}$ is able to distinguish between the clean and adversarial images. The output logit $\mathbf{w}$ of $f_{p}$ represents the probability of an image to be clean or adversarial. Therefore, we use $\mathbf{w}$ to ensemble the target model and the defense model:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (13)
In this way, both the clean images and the adversarial images are correctly classified.
B Premium-CAD
We propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In each stage, we train an attack-specific defense model using abundant feedback and ensemble the target model and all the defense models by a stage-evolved weight estimator.
Attack-Specific Defense Model
In stage 00, the initial defense model is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. In stage $i$, another defense model is trained using abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)\}_{N\times K}$ where $N$ is the number of classes in dataset $\mathcal{D}_{\mathrm{train}}$ and $K$ is the amount of feedback per class. The scale of feedback is set to $K=10$ in CAD and $K=1000$ in premium-CAD.
In this way, we have $i+1$ defense models in stage $i$. The defense model is designed to tackle adversarial examples from a specific attack, complementing each other.
Algorithm 3 Premium Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:   for $i$ in $1,\cdots,T$ do
2:      Train an attack-specific defense model $f_{i}$.
3:      Train a stage-evolved weight estimator $f_{p}^{i}$.
4:      Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Eq. 13.
5:   end for
Stage-Evolved Weight Estimator
To ensemble the target model $f_{t}$ and defense models $[f_{0},f_{1},\cdots,f_{i}]$, we use a joint dataset $\mathcal{D}_{J}^{i}=\bigcup_{i=0,1,\cdots,i}\mathcal{A}_{\mathrm{train}}^{i}\cup\mathcal{D}_{\mathrm{train}}$ to train the weight estimator $f_{p}^{i}$ in stage $i$. Feed an image to the weight estimator $f_{p}^{i}$, it output a weight vector $\mathbf{w}^{i}\in\mathbb{R}^{i+2}$ to ensemble models:
$\mathbf{w}^{i}=f_{i}(\mathbf{x})$ (14)
The final output for premium-CAD is:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}^{i}\cdot[f_{t}(\mathbf{x}),f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\cdots,f_{i}(\mathbf{x})]^{\top}$ (15)
The overall algorithm of premium-CAD framework is presented in Algorithm 3. The loss function for training attack-specific defense models and stage-evolved weight estimators is Cross-Entropy Loss.

### v2

@@
-Continual Adversarial Defense
+
 Abstract
-In response to the rapidly evolving nature of adversarial attacks on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible.
-However, designing a defense method that can generalize to all types of attacks, including unseen ones, is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks used by many attackers.
-The defense system needs to upgrade itself by utilizing few-shot defense feedback and efficient memory.
+In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible.
+However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on.
+The defense system must gather online few-shot defense feedback to promptly enhance itself, leveraging efficient memory utilization.
 Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage.
 In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images.
-We leverage cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles.
-Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of 10 modern adversarial attacks and significant improvements over 10 baseline methods.
-In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against old attacks.
+We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles.
+Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods.
+In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against previous attacks.
 Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
-[^1] Code at https://github.com/cc13qq/CAD
-Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender.
-During the $i$-th stage, the defender $f_{i}$ leverages few-shot defense feedback provided by the security department and users of the target model to adapt to new attacks.
-The defender also needs to efficiently use memory to store the cache, retaining knowledge of past attacks.
+Keywords:
+Adversarial Defense Continual Learning
+Figure 1: In a dynamic environment, the target model $f_{t}$ encounters a wide spectrum of attacks across various stages while being safeguarded by the defender.
+During the $i$-th stage, the defender leverages few-shot defense feedback, which is collected by the online defense system or provided by the security department and users of the target model, to adapt to emerging threats.
+Furthermore, the defender is required to efficiently manage memory utilization to store the cache, preserving insights gained from past attacks.
 1 Introduction
@@
 From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones.
-The truth is, the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32].
+The truth is that the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32].
 As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [20].
 The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [13].
-To simulate a dynamic environment more realistically, we propose a defense scenario that considers practicality.
-As illustrated in Fig.
-1, the attacker employs diverse attacks at different stages to target the model, while the defender adapts to emerging attacks by leveraging few-shot defense feedback.
-In this paper, we use adversarial examples with ground truths as the defense feedback, which is provided by the security department and users of the target model in real-world situations.
-In this ongoing “arms race,” the attacker operates within the gray-box setting [32], where they possess knowledge of the classifier’s architecture and have access to training data but are unaware of the defense mechanism in place.
-On the other hand, the defender has the same training data and access to the target model, but also possesses full knowledge of the initial attack.
+Inspired by online learning, the crucial approach to addressing dynamic scenarios lies in embracing a perpetually evolving defense strategy.
+Consequently, deploying a defense system on the cloud that continuously gathers attack examples from newly emerged adversarial algorithms and promptly adapts to these attacks represents a promising method for responding to such scenarios.
+As depicted in Fig.
+1, a variety of attacks are detected across different stages, with the defender adapting to emerging threats by utilizing few-shot defense feedback.
+In this paper, we employ adversarial examples with ground truths as the defense feedback, sourced either from the online defense system’s crawling efforts or provided by the security department and users of the target model in real-world scenarios.
+The defender undergoes training with an initial attack and shares the same training data and architecture as the target model.
+Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
 Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting.
-As defenders, it’s essential to adapt to various new attacks at different stages while retaining knowledge of previous ones.
+As defenders, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters.
 (2) Few-shot adaptation.
-Increased feedback indicates more successful attacks on the target model.
-Therefore, we cannot afford too many successful attacks before upgrading the defense.
+The escalation in feedback signifies a greater potential for attacks on the target model.
+Therefore, we cannot tolerate an excessive number of attacks before adaptation.
 (3) Memory-efficient adaptation.
-Over time, a continuous influx of attacks results in accumulating defense feedback, potentially leading to memory constraints.
-In practical terms, we may not have sufficient memory capacity to withstand this.
+Over time, a continuous influx of attacks leads to the accumulation of defense feedback, potentially posing memory constraints.
+In practical terms, we may not have sufficient memory capacity to withstand this pressure.
 (4) High accuracy in classifying both clean and adversarial images.
 A robust defense should not compromise the interests of those it protects.
-Hence, maintaining high classification accuracy on both clean and adversarial images is crucial.
+Hence, ensuring high classification accuracy on both clean and adversarial images is paramount.
 In this paper, we propose the Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense feedback and efficient memory.
@@
 Our main contributions can be summarized as follows:
-$\bullet$ In order to simulate the rapidly evolving adversarial attacks against image classifiers, we introduce, for the first time, a dynamic scenario of various attacks that emerge in different stages.
-The defender is required to adapt to new attacks using few-shot defense feedback and efficient memory, while also retaining knowledge of previous attacks and upholding high performance on both clean and adversarial images.
+$\bullet$ To withstand the rapidly evolving adversarial attacks targeting image classifiers, we introduce, for the first time, a dynamic scenario featuring diverse attacks emerging at different stages to continuously enhance the defense.
+The defender must adeptly adapt to novel attacks using few-shot defense feedback and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial images.
 $\bullet$ We propose the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial images.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
-$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory, and demonstrate significant improvements over 10 baseline methods.
+$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory and demonstrate significant improvements over 10 baseline methods.
 2 Related Work
@@
 [26] propose DiffPure that uses the forward and reverse processes of diffusion models to purify adversarial images.
-Certainly, both adversarial training and purification methods have shown limitations in robustness, leading to reduced classification performance on clean images.
-Furthermore, the dynamic nature of the environment introduces a diverse array of unique attacks employed by multiple attackers.
-In this paper, we propose Continual Adversarial Defense (CAD) framework as a solution to effectively combat various attacks emerging stage by stage, while maintaining strong performance in classifying clean images.
+Indeed, both adversarial training and purification methods have exhibited limitations in robustness, resulting in diminished classification performance on clean images.
+Moreover, the dynamic environment introduces a multitude of unique attacks, rendering a single defense inadequate.
+In this paper, we present the Continual Adversarial Defense (CAD) framework as a solution to effectively counter various attacks emerging stage by stage, while ensuring robust performance in classifying clean images.
 2.3 Continual Learning
-Continual learning (CL) [21, 28, 19, 18, 20, 25, 22, 24, 46, 23] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks.
+Continual learning (CL) [22, 27] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks.
 Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [27], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [47].
@@
 3 Threat Model and Defense Principles
-The defense system operates within a dynamic environment characterized by a multitude of unique attacks since the growing number of defenses naturally encourages the invention of new attack algorithms.
-Once frustrated by an existing defense, attackers undoubtedly resort to deploying various attacks in pursuit of breakthroughs.
-In response to this reality, we simulate the dynamic environment and put forth a practical threat model.
-In this threat model, attackers deploy different attacks in various stages to assault the target model, and the defender must adapt to evolving attacks by utilizing the defense feedback, just like an ”arms race”.
-The defense feedback is typically provided by security services and users of the target model upon the discovery of successful attacks during screening or cleanup efforts in the real world, and plays a critical role in informing and shaping the defender’s response to emerging threats.
+The defense system is deployed online and functions within a dynamic environment marked by a plethora of unique attacks, driven by the proliferation of defenses that naturally prompt the development of new attack algorithms.
+As time progresses, numerous new attacks, accompanied by attack examples, will emerge on the internet, aiming to exploit vulnerabilities in the defense system and achieve breakthroughs.
+In response to this reality, we simulate the dynamic environment and propose a practical threat model.
+In this model, distinct attacks are identified across various stages, requiring the defender to adapt to evolving threats through the utilization of defense feedback.
+Typically, defense feedback is obtained through crawling efforts of the online defense system or is provided by security services and users of the target model.
+This feedback is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world.
 In this paper, we use adversarial examples with ground truths as the defense feedback.
 We formulate this attack vs defense scenario as follows.
-Our scenario is established on the $N$-way $K$-shot classification paradigm.
-The attacker attacks under the gray-box setting [32] in which the attacker knows the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space.
-The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage.
-At the $i$-th stage, where $i=1,2,\cdots,T$, the attacker knows the failure of previous attacks and turns to a new attack $A_{i}(\cdot)$.
-Meanwhile, the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack.
-Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
 Figure 2: Illustration of CAD.
@@
 In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning.
-And then, a weight estimator model $f_{p}$ is trained for the model ensemble.
+Then, a weight estimator model $f_{p}$ is trained for model ensemble.
 In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback.
 To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process.
-After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator to maintain high accuracy of both clean and adversarial images.
+After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high accuracy in classifying clean and adversarial images.
+Our scenario is established on the $N$-way $K$-shot classification paradigm.
+Attacks are launched under the gray-box setting [32] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space.
+The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage.
+At the $i$-th stage, where $i=1,2,\cdots,T$, a new attack $A_{i}(\cdot)$.
+is detected and the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack.
+Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
 Taking reality into account, the defense mechanism should satisfy the following principles:
@@
 Continual adaptation to new attacks without catastrophic forgetting.
-In the dynamic environment, it is not uncommon for certain attackers to employ the same attack across different stages.
-Therefore, the defense mechanism must be capable of adapting to a range of new attacks in each stage while also retaining knowledge of previous ones.
+In the dynamic environment, various new attacks will emerge on the internet as time goes on.
+Therefore, the defense mechanism must be capable of adapting to a range of new attacks across different time stages while also retaining knowledge of previous ones.
 Principle 2
@@
 Memory-efficient adaptation.
-Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to the memory constraints.
+Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to memory constraints.
 In practical terms, the defender may not have sufficient memory space to accommodate this accumulation.
@@
 The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation.
-This decline could potentially have a negative impact on crucial real-world business operations.
-As a result, it is imperative for a defender to prioritize maintaining performance on both clean and adversarial images.
+This decline could potentially hurt crucial real-world business operations.
+As a result, a defender must prioritize maintaining performance on both clean and adversarial images.
 4 Continual Adversarial Defense Framework
@@
 In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$.
-Inspired by continual-learning (CL) [28], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
-Following [21], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ at the $i$-th stage.
+Inspired by continual learning (CL) [27], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
+Following [22], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage.
 The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion.
@@
 To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
-First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ‘virtual classes [43], where $V=N\times T$ is the number of virtual classes, i.e.
+First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as "virtual classes" [43], where $V=N\times T$ is the number of virtual classes, i.e.
 the reserved classes for future attacks.
@@
 4.3 Prototype Augmentation for Memory Efficiency
-In response to Principle 3, we use prototype augmentation to efficiently use memory.
+In response to Principle 3, we use prototype augmentation to efficiently manage memory.
 When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50].
-Many CIL methods store a fraction of old data to jointly train the model with current data to address this issue [28].
+Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [27].
 However, preserving old feedback may also lead to memory shortage.
@@
 $\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{proto}}$ (9)
+4.4 Model Ensemble for Clean and Adversarial Image Classification
+To maintain high classification accuracy on both clean and adversarial images mentioned in Principle 4, we propose the model ensemble as the last part of CAD.
+Ensemble adversarial training (EAT) [34] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean images.
+We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models
+We adopt the self-perturbation [37] to train the weight estimator at the first stage.
+Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial.
+Details of the self-perturbation are given in the supplementary.
+The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
+$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (10)
+In this way, both the clean images and the adversarial images are correctly classified.
+The overall algorithm of CAD framework is presented in Algorithm 1.
 Algorithm 1 Continual Adversarial Defense
-Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
+Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
 Output: Defense model ensemble.
@@
 10:  end for
-4.4 Model Ensemble for Clean and Adversarial Image Classification
-To maintain high classification accuracy on both clean and adversarial images mentioned in Principle 4, we propose the model ensemble as the last part of CAD.
-Ensemble adversarial training (EAT) [34] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean images.
-We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models
-We adopt the self-perturbation [37] to train the weight estimator at the first stage.
-Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial.
-Details of the self-perturbation are given in the supplementary.
-The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
-$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (10)
-In this way, both the clean images and the adversarial images are correctly classified.
-The overall algorithm of CAD framework is presented in Algorithm 1.
 5 Experiment
-To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from two research streams.
-For evaluation and analysis, we use two metrics in this section: (1) classification accuracy against each attack after adaptation.
-(2) the average of the accuracy of all attacks that have occurred in and before in each stage, i.e.
-average incremental accuracy [28].
+To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from three research streams.
+For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy against each attack after adaptation for defense performance evaluation.
+(2) Average accuracy of all attacks occurring within and before each stage, i.e.
+average incremental accuracy [27], for assessing the knowledge retention of the model.
 Datasets
@@
 Attack Algorithms
-We select 11 adversarial attack methods PGD [26], FGSM [6], BIM [14], RFGSM [35], MIM [3], DIM [40], NIM [17], SNIM [17], VNIM [38], and VMIM [38] to compose the attack pool.
-The perturbation magnitude of all the attacks is set to $\epsilon=8$ under $l_{\infty}$ following [44].
+We choose PGD-$l_{\infty}$ [23] as the initial attack.
+After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM [18], RFGSM [34], MIM [6], DIM [40], NIM [21], SNIM [21], VNIM [38] and VMIM [38] to compose the attack pool in defense performance evaluation.
+The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44].
+As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM [38] and SNIM [21], $l_{2}$ attack CW [1] and DeepFool [25], $l_{1}$ attack EAD [2] and EADEN [2], and $l_{0}$ attack OnePixel [31] and SparseFool [24] compose the attack pool.
+Details of attack parameters are given in the supplementary.
+Stage 0 1 2 3 4 5 6 7 8 -
+Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
+None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
+TRADES [44] 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
+JEM [10] 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
+PAT [19] 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
+GAIRAT [45] 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
+FastAdv [14] 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
+RPF [5] 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
+DMAT [39] 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
+EBM [12] 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
+ADP [41] 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
+DiffPure [26] 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
+CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
+Table 1: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10.
+Volumn ”Clean” represents the standard accuracy of clean images.
+$\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense feedback.
 Implementation Details
@@
 The amount of feedback per class is set to $K=10$.
-The number of stages is set to $T=9$ since there are 10 attacks (PGD is the initial attack).
+The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack).
 Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [47].
 Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
-Baselines
-We first compare to adversarial training methods TRADES [44], JEM [7], GAIRAT [45], FastAdv [11], RPF [2] and DMAT [39].
+Defense Baselines
+We first compare to adversarial training methods TRADES [44], JEM [10], GAIRAT [45], FastAdv [14], RPF [5] and DMAT [39], and purification methods EBM [12], ADP [41] and DiffPure [26] in defense performance evaluation.
+As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC [16], SSRE [51], BEEF [36], and FACT [47].
 Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback.
-In premium-CAD, we train additional attack-specific defense models using abundant feedback (i.e., 1000 images per class ) in each stage, and ensemble them with the target model by the weight estimator.
-Details about premium-CAD are given in the supplementary.
-Stage 0 1 2 3 4 5 6 7 8 9 -
-Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
-TRADES [44] 55.8 57.4 60.7 55.9 55.8 57.1 57.1 58.9 60.2 57.3 84.8
-JEM [7] 55.8 46.5 60.7 52.3 52.4 51.1 51.4 57.0 62.0 48.2 91.7
-PAT [15] 45.5 43.8 49.8 45.5 45.3 45.2 45.2 43.9 45.1 45.4 69.3
-GAIRAT [45] 66.9 66.4 68.3 66.9 66.7 66.5 66.5 68.6 70.4 66.4 89.4
-FastAdv [11] 37.1 36.8 41.9 37.0 37.7 37.7 54.6 36.4 39.6 37.6 75.2
-RPF [2] 60.9 52.2 58.0 58.9 54.9 57.6 57.5 53.5 52.8 52.6 83.5
-DMAT [39] 71.7 73.1 75.8 71.7 71.7 73.0 73.0 75.5 76.0 73.0 92.4
-EBM [9] 75.2 73.0 70.0 76.6 76.5 72.4 73.9 74.0 72.9 73.8 86.8
-ADP [41] 85.4 79.3 82.0 84.4 85.1 84.1 83.7 81.0 80.5 79.1 80.5
-DiffPure [27] 72.5 69.8 71.8 70.1 72.2 73.4 73.5 72.7 70.2 69.4 89.3
-CAD(ours) 95.8 93.4 83.8 94.7 95.0 93.7 93.7 95.2 95.4 93.7 95.8
-$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.6 94.6 94.1 94.6 94.3 95.1 95.1 95.5 95.7 96.0
-Table 1: Classification accuracy (%) against various attacks on CIFAR-10.
-Volumn ”Clean” represents the standard accuracy of clean images.
-$\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by the efficient memory and few-shot defense feedback.
-Stage 0 1 2 3 4 5 6 7 8 9 -
-Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
-TRADES [44] 46.5 49.6 48.8 46.2 48.0 47.7 46.7 48.2 45.5 49.5 76.2
-GAIRAT [45] 59.8 58.4 59.4 59.2 58.8 58.9 58.9 59.7 60.2 59.8 74.1
-RPF [2] 56.4 51.3 55.8 56.6 54.4 52.6 53.9 53.3 49.9 50.1 69.4
-DMAT [39] 64.3 66.2 66.4 65.9 65.4 64.4 64.1 65.8 65.1 65.6 75.9
-ADP [41] 77.2 71.8 73.7 71.8 75.6 75.9 76.2 76.0 76.9 72.3 74.1
-DiffPure [27] 53.2 52.4 53.7 55.8 53.9 53.4 52.6 54.7 53.0 52.3 73.6
-CAD(ours) 83.7 82.0 76.2 82.5 83.5 82.0 81.9 82.2 83.4 81.9 88.1
-$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 84.0 84.1 82.6 84.3 83.8 84.2 82.2 83.9 81.4 87.8
-Table 2: Classification accuracy (%) on ImageNet-100.
+In premium-CAD, we train additional attack-specific defense models using abundant feedback in each stage, and ensemble them with the target model by the weight estimator.
+Details about the implementation of defense baseline and premium-CAD are given in the supplementary.
 5.1 Comparisons to Baselines
-We compared CAD to other baseline methods in defending against 10 adversarial attacks.
-In our approach, the defense environment is dynamic, with attacks occurring at different stages.
-In contrast, the other baseline methods operate in a static defense environment that does not differentiate the order of attacks.
-We report the robust accuracy, which measures the performance on adversarial examples generated by attacks, and the standard accuracy, which measures the performance on clean data, separately for each method.
-We compare the defense performance of our method to other methods on CIFAR-10 and ImageNet-100.
-All the evaluation is conducted on the whole test set.
+To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks.
+For CAD, the defense environment is dynamic, with attacks occurring at different stages.
+The other baseline methods operate in a static defense environment that does not differentiate the order of attacks.
+All the evaluation is conducted on the whole test set of CIFAR-10 and ImageNet-100.
+Classification accuracy is reported in Tab.
+1 and Tab.
+2.
+Stage 0 1 2 3 4 5 6 7 8 -
+Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
+None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
+TRADES [44] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
+GAIRAT [45] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
+RPF [5] 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
+DMAT [39] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
+ADP [41] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
+DiffPure [26] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
+CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
+Table 2: Classification accuracy(%) against various attacks($l_{\infty}$) on ImageNet-100.
+Stage 0 1 2 3 4 5 6 7 8
+Attack PGD SNIM VNIM CW DeepF. EAD EADEN OnePixel SparseF.
+None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
+EWC [16] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
+SSRE [51] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
+BEEF [36] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
+FACT [47] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
+CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
+Table 3: Continual defense performance (Average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10.
+DeepF.
+and SparseF.
+denote DeepFool [25] and SparseFool [24] respectively.
 As depicted in Tab.
 1 and Tab.
-2, despite the modest scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack.
-When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ in CIFAR-10 and $6\%$ in ImageNet-100 for each attack.
+2, despite the small scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack.
+When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ on CIFAR-10 and $6\%$ on ImageNet-100 for each attack.
 This underscores its capacity for continuous adaptation in response to emerging attacks.
-Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy for CIFAR-10 and $88.1\%$ for ImageNet-100, while the testing accuracy of the target model stands at $96.4\%$ for CIFAR-10 and $89.8\%$ for ImageNet-100.
+Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy on CIFAR-10 and $88.1\%$ on ImageNet-100, while the target model stands at $96.4\%$ on CIFAR-10 and $89.8\%$ on ImageNet-100.
 This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
-We observe that the accuracy against FGSM is relatively lower than that for other attacks, largely due to the distinct noise pattern of the one-step FGSM compared to other attacks.
-To assess CAD’s performance in mitigating catastrophic forgetting, we present the average incremental accuracy in Fig.
-3.
-The curve initially experiences a slight decline in the first three stages before stabilizing.
-This temporary dip is primarily attributable to FGSM’s influence, which pulls down the entire curve, and the adaptation to new attacks, resulting in a minor decrease in accuracy against previous attacks.
-Subsequently, the curve exhibits a modest increase due to the resemblance of other attacks to the initial attack PGD, thereby elevating the overall curve.
-Throughout the evaluation, the curves consistently hover around $92\%$ for CIFAR-10 and $80\%$ for ImageNet-10, underscoring CAD’s robust capacity to counter catastrophic forgetting.
-Note that premium-CAD outperforms all baseline methods. Given adequate defense feedback and memory resources, premium-CAD emerges as the superior choice.
+To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods through experiments on CIFAR-10 with attacks under different norms $l_{p=0,1,2,\infty}$.
+All baseline methods are functioned within our dynamic scenario.
+As depicted in Tab.
+3, both CAD and premium-CAD maintain robust defense performance despite changes in attack types.
+However, as the variety of attack types increases, the defense performance gradually diminishes.
+This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense.
+Compared with the top baseline method FACT [47], CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
+It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense feedback and memory resources, premium-CAD emerges as the superior choice.
+(a) PGD
+(b) CW
+(c) EAD
+(d) OnePixel
+Figure 3: T-SNE diagrams of features extracted from adversarial test data encoded by the defense model after all stages.
+Each color represents a class of CIFAR10.
+After several steps of adaptation, CAD retains the capability to categorize the attacked data.
 5.2 Ablation Study
-Figure 3: Average incremental accuracy (%) of CAD.
-As argued, embedding reservation allows the defense model to be adapted more easily in the future and alleviates overfitting, prototype augmentation maintains the decision boundary of previous stages, and model ensemble ensures robust classification performance for both clean and adversarial images.
-Subsequently, we proceed to undertake an ablation study to validate the efficacy of each component.
+As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting.
+Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
+Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and
+Stage 0 1 2 3 -
+Attack PGD CW EAD OnePixel Clean
+w/o $L_{\mathrm{virtual}}$ 94.5 82.4 76.1 71.8 95.8
+w/o $L_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
+w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
+CAD(ours) 94.5 89.9 86.6 85.3 95.8
+Table 4: Average incremental accuracy(%) on CIFAR-10 for ablation.
+OnePixel($l_{0}$) to validate the efficacy of each component.
 As shown in Tab.
-3 and Fig.
-4 (Left), CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.2\%$ and $2.2\%$ respectivily in performance for adaptation.
-Furthermore, the absence of model ensemble results in the defense model’s inability to uphold classification accuracy for clean images alone.
-Additionally, we present curves illustrating the classification accuracy of each attack in Fig.
-4 (Right).
-These curves demonstrate that the adaptation of CAD to FGSM is comparatively less effective than for other attacks, aligning with our earlier observations about FGSM in Sec.
-5.1.
-Stage 0 1 2 3 -
-Attack PGD VNIM FGSM BIM Clean
-w/o $L_{\mathrm{virtual}}$ 95.6 86.6 70.6 78.5 95.8
-w/o $L_{\mathrm{proto}}$ 95.7 93.7 81.6 88.2 95.8
-w/o $f_{p}$ 95.8 93.4 83.8 94.7 55.9
-CAD(ours) 95.8 93.4 83.8 94.7 95.8
-Table 3: Classification accuracy (%) on CIFAR-10 for ablation.
-Figure 4: Accuracy (%) in each stage on CIFAR-10. Left: average incremental accuracy of ablation. Right: classification accuracy against each attack.
+4, CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage.
+The absence of model ensemble leads to the inability to uphold performance on clean images.
+Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Fig.
+3 to visualize the defense capability of CAD after adaptation.
+These features are encoded by the defense model after the last stage of adapting to OnePixel attack.
+From (a) to (d) in Fig.
+3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types.
+After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
 5.3 Multiple Attacks in One Stage
-In practice, it is common for multiple attacks to occur simultaneously within a single stage.
-To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks.
+In practice, it is common for multiple attacks to occur simultaneously within a
+Stage 0 1 2 3
+$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
+$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
+$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
+Table 5: Classification accuracy(%) about the number of attacks($l_{\infty}$) in one stage.
+single stage.
+To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks on CIFAR-10.
 As presented in Tab.
-4, when the number of attack increase to $2$ and $3$, the classification accuracy against these attacks still remain at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively.
-The results presented in Tab.
-4 indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
+5, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively.
+The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
+5.4 Number of Shots for Adaptation
+As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we
 Stage 0 1 2 3
-$\mathrm{number}=1$ 95.8 93.4 83.8 94.7
-$\mathrm{number}=2$ 95.8 88.5 93.5 92.8
-$\mathrm{number}=3$ 95.8 89.5 93.6 92.5
-Table 4: Classification accuracy (%) on CIFAR-10 for different number of attacks in one stage.
-5.4 Number of Shots for Adaptation
-As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback.
-To investigate the influence of the amount of feedback $K$ per class, we conducted tests with $K=15,10,5,1$.
+Attack PGD CW EAD OnePixel
+$K=15$ 94.5 90.0 88.4 85.5
+$K=10$ 94.5 89.9 86.6 85.3
+$K=5$ 94.5 89.4 86.1 84.6
+$K=1$ 94.5 87.9 84.6 83.4
+Table 6: Average incremental accuracy(%) with different feedback scale.
+conducted tests with $K=15,10,5,1$ on CIFAR-10.
 The results presented in Tab.
 6 illustrate that defense performance diminishes as $K$ decreases.
-Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($90.2\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
-Stage 0 1 2 3
-Attack PGD VNIM FGSM BIM
-$K=15$ 95.8 93.7 81.2 94.8
-$K=10$ 95.8 93.4 83.8 94.6
-$K=5$ 95.8 93.2 80.8 93.7
-$K=1$ 95.8 92.8 77.5 90.2
-Table 5: Classification accuracy (%) on CIFAR-10 with different feedback scale $K$.
+Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($93.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
 5.5 Memory Allowance for Adaptation
 As outlined in Principle 3, we need to employ memory efficiently for continual defense.
-In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each stage.
-As depicted in Tab.
-6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size.
-This underscores our approach’s ability to utilize efficient memory.
+In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
 Cache Size Model Size
@@
 $\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
-Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD on CIFAR-10.
+Table 7: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
+stage on CIFAR-10.
+As depicted in Tab.
+7, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size.
+This underscores our approach’s ability to utilize efficient memory.
 6 Conclusion
-The operational environment for defense systems is inherently dynamic, and it is unrealistic to expect a defense method to address all types of attacks.
-As a response to this situation, we propose the first continual adversarial defense (CAD) framework, designed to dynamically adapt to a wide array of attacks as they emerge in sequential stages.
+The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks.
+In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework.
+CAD is engineered to collect few-shot defense feedback from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages.
 In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles.
 Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
-7 Acknowledgements
-This work was supported in part by the Natural Science Foundation of China under Grant 61972169,62372203 and 62302186, in part by the National key research and development program of China(2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042), in part by the Knowledge Innovation Program of Wuhan-Basic Research, in part by China Postdoctoral Science Foundation 2022M711251.
-Supplementary Material
 A Weight Estimator Base on Self-Perturbation
-In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation to train a weight estimator to ensemble the target model and the defense model.
+In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation [37] to train a weight estimator to ensemble the target model and the defense model.
 We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples.
@@
 The overall algorithm of premium-CAD framework is presented in Algorithm 3. The loss function for training attack-specific defense models and stage-evolved weight estimators is Cross-Entropy Loss.
+C Implementation Details
+C.1 Adversarial Attack
+In this work, all of the adversarial attacks are implemented using TorchAttacks [15].
+The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44].
+Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively.
+The number of perturbed pixels of OnePixel [31] is set to $50$.
+The other parameters are the default ones in TorchAttacks.
+C.2 Defense Baselines
+The adversarial training baselines are implemented using RobustBench [3], while purification baselines are implemented using open-source code and model checkpoints.
+CL baselines are implemented using PyCIL [48].
+The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100.
+CL models are trained under the same settings as CAD.
+Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.

### v3

@@
-
+Continual Adversarial Defense
 Abstract
@@
 However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on.
-The defense system must gather online few-shot defense feedback to promptly enhance itself, leveraging efficient memory utilization.
-Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage.
-In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images.
+A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself.
+Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data.
 We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles.
-Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods.
-In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against previous attacks.
+Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods.
+In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks.
 Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
-Keywords:
-Adversarial Defense Continual Learning
-Figure 1: In a dynamic environment, the target model $f_{t}$ encounters a wide spectrum of attacks across various stages while being safeguarded by the defender.
-During the $i$-th stage, the defender leverages few-shot defense feedback, which is collected by the online defense system or provided by the security department and users of the target model, to adapt to emerging threats.
-Furthermore, the defender is required to efficiently manage memory utilization to store the cache, preserving insights gained from past attacks.
-1 Introduction
-In recent years, deep neural networks have been widely used in numerous vision tasks [29], leading to remarkable breakthroughs.
-However, many neural networks are vulnerable to adversarial attacks, which aim to deceive a classifier by adding subtle perturbations to input images, thus altering prediction results.
-This vulnerability seriously jeopardizes the reliability of deep neural networks, particularly in security- and trust-sensitive domains.
-Researchers have proposed various general defense methods against adversarial attacks.
-As a standard defense method, adversarial training [44] aims to enhance the robustness of the target model by training it with adversarial examples.
-Another branch of adversarial defense involves purifying the data stream [41] to remove potential adversarial perturbations or noise that could deceive the model.
-However, models of adversarial training often exhibit reduced classification accuracy on adversarial examples and may sacrifice classification capacity on clean images.
-On the other hand, purification techniques may appear more reliable, but the denoising procedure can inadvertently smooth the high-frequency texture of clean images, potentially causing issues for the target model.
-From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones.
-The truth is that the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32].
-As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [20].
-The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [13].
-Inspired by online learning, the crucial approach to addressing dynamic scenarios lies in embracing a perpetually evolving defense strategy.
-Consequently, deploying a defense system on the cloud that continuously gathers attack examples from newly emerged adversarial algorithms and promptly adapts to these attacks represents a promising method for responding to such scenarios.
-As depicted in Fig.
-1, a variety of attacks are detected across different stages, with the defender adapting to emerging threats by utilizing few-shot defense feedback.
-In this paper, we employ adversarial examples with ground truths as the defense feedback, sourced either from the online defense system’s crawling efforts or provided by the security department and users of the target model in real-world scenarios.
-The defender undergoes training with an initial attack and shares the same training data and architecture as the target model.
+Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender.
+During the $i$-th stage, the defense model leverages few-shot defense budget provided by the web crawler, security department, and users of the target model to adapt to new attacks.
+The defense model also needs to efficiently use memory to store the cache while retaining knowledge of past attacks.
+Introduction
+Adversarial attack (Madry et al.
+2017) aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust.
+In order to promote the robustness of DNN, adversarial training (Zhang et al.
+2019) and purification techniques (Yoon, Hwang, and Lee 2021) are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage (Zhou and Hua 2024), but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
+Contrary to the one-shot defense assumption, the defense system deployed in real scenarios encounters a dynamic environment characterized by numerous unique attacks with higher aggressivity (Li et al.
+2020).
+A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms.
+A concurrent continual defense work (Zhou and Hua 2024) treats attacks as a task sequence and tries to train a robust model through lifelong learning.
+However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
+To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model.
+As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e.
+adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios.
+In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget.
+When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory.
+The defense system also has access to an initial attack and shares the same training data and architecture as the target model.
 Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
-Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting.
-As defenders, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters.
+Taking real-world considerations into account, we propose four principles for continual defense: (1) Continual adaptation to new attacks without catastrophic forgetting.
+As a defender, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters.
 (2) Few-shot adaptation.
-The escalation in feedback signifies a greater potential for attacks on the target model.
+The escalation in the defense budget signifies a greater potential for attacks on the target model.
 Therefore, we cannot tolerate an excessive number of attacks before adaptation.
 (3) Memory-efficient adaptation.
-Over time, a continuous influx of attacks leads to the accumulation of defense feedback, potentially posing memory constraints.
+Over time, a continuous influx of attacks leads to the accumulation of defense budget, potentially posing memory constraints.
 In practical terms, we may not have sufficient memory capacity to withstand this pressure.
-(4) High accuracy in classifying both clean and adversarial images.
+(4) High accuracy in classifying clean and adversarial data.
 A robust defense should not compromise the interests of those it protects.
-Hence, ensuring high classification accuracy on both clean and adversarial images is paramount.
-In this paper, we propose the Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense feedback and efficient memory.
-In the first stage, we train an initial defense model using adversarial images from the initial attack.
-This model is specifically designed to classify images with adversarial noise, complementing the target model.
-Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes.
-We then fine-tune this expanded layer using few-shot defense feedback.
-To address the over-fitting issue that arises from the few-shot fine-tuning process, we reserve embedding space in the defense model for future attacks.
-This is achieved by generating and assigning virtual prototypes, which help compress the embedding of previous attacks.
-To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any feedback explicitly.
-Simultaneously, we utilize a small model to ensemble the defense model with the target model by estimating reliable logits for input images, ensuring maximum classification accuracy for both clean and adversarial images.
-Based on our comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot feedback while maintaining high accuracy on clean images.
+Hence, ensuring high classification accuracy on both clean and adversarial data is paramount.
+In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory.
+In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model.
+Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget.
+To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks.
+To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly.
+After adaptation of one stage, we refine the classifier to prevent undesirable expansion.
+Simultaneously, a small model is utilized to ensemble the defense model with the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
+Based on comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot defense budget while maintaining high accuracy on clean data.
+To the best of our knowledge, this is the first time to discuss the concept and deployment scenario of continual adversarial defense.
 Our main contributions can be summarized as follows:
-$\bullet$ To withstand the rapidly evolving adversarial attacks targeting image classifiers, we introduce, for the first time, a dynamic scenario featuring diverse attacks emerging at different stages to continuously enhance the defense.
-The defender must adeptly adapt to novel attacks using few-shot defense feedback and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial images.
-$\bullet$ We propose the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial images.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
-$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory and demonstrate significant improvements over 10 baseline methods.
-2 Related Work
-2.1 Adversarial Attacks
-Adversarial attacks [9] attempt to fool a classifier into changing prediction results by attaching subtle perturbations to the input images while maintaining imperceptibility from human eyes.
-PGD [23] and BIM [18] generate adversarial examples through several iterations of perturbation and get significantly improved attack performance.
-DIM [40], TIM [7], and SIM [21] improve the transferability under the black-box setting and breach several defense techniques.
-Incorporating variance tuning, VIM [38] achieves a high attack rate against multiple advanced defense methods.
-In this paper, a wide spectrum of attacks is used to offense the target model at different stages under the gray-box setting.
-2.2 Adversarial Defense
-Adversarial training [23, 45] is the mainstream of adversarial defense that uses attacked samples to train a robust model.
-TRADES [44] characterize the trade-off between accuracy and robustness for classification problems via decomposing the robust error as the sum of the natural error and the boundary error.
-JEM [10] uses energy-based training of the joint distribution and improves the calibration and robustness of target models.
-Replacing the denoising diffusion probabilistic model with the elucidating diffusion model, DMAT [39] improves adversarial training and achieves new SOTA robust accuracy.
-Another technical solution for protecting DNNs from adversarial attacks is adversarial purification [28, 12], which aims to purify the data stream before feeding it into the target model, functioning under the gray-box setting [32].
-Based on an EBM trained with Denoising Score-Matching, Yoon et al.
-[41] propose Adaptive Denoising Purification (ADP) which conducts purification in a randomized purification scheme.
-Nie et al.
-[26] propose DiffPure that uses the forward and reverse processes of diffusion models to purify adversarial images.
-Indeed, both adversarial training and purification methods have exhibited limitations in robustness, resulting in diminished classification performance on clean images.
-Moreover, the dynamic environment introduces a multitude of unique attacks, rendering a single defense inadequate.
-In this paper, we present the Continual Adversarial Defense (CAD) framework as a solution to effectively counter various attacks emerging stage by stage, while ensuring robust performance in classifying clean images.
-2.3 Continual Learning
-Continual learning (CL) [22, 27] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks.
-Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [27], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [47].
-In recent years, some methods have aimed to solve the CL problem without depending on preserving data (called none-exemplar) [51], and some methods have attempted to learn new classes in a few-shot scenario [43] in which only a small number of new class data is gainable.
-In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense feedback and efficient memory for adaptation.
-3 Threat Model and Defense Principles
-The defense system is deployed online and functions within a dynamic environment marked by a plethora of unique attacks, driven by the proliferation of defenses that naturally prompt the development of new attack algorithms.
-As time progresses, numerous new attacks, accompanied by attack examples, will emerge on the internet, aiming to exploit vulnerabilities in the defense system and achieve breakthroughs.
-In response to this reality, we simulate the dynamic environment and propose a practical threat model.
-In this model, distinct attacks are identified across various stages, requiring the defender to adapt to evolving threats through the utilization of defense feedback.
-Typically, defense feedback is obtained through crawling efforts of the online defense system or is provided by security services and users of the target model.
-This feedback is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world.
-In this paper, we use adversarial examples with ground truths as the defense feedback.
-We formulate this attack vs defense scenario as follows.
+$\bullet$ To withstand the rapidly evolving adversarial attacks targeting DNNs, we put forward the deployment of continual defense strategy and propose a practical threat model in which diverse attacks emerge at different stages.
+The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial data.
+$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
+$\bullet$ Extensive experiments validate the effectiveness of CAD against multiple stages of modern adversarial attacks using few-shot budget and efficient memory and demonstrate significant improvements over baseline methods.
+Related Work
+Adversarial Defense
+Researchers have proposed various general defense methods against adversarial attacks (Madry et al.
+2017; Wu et al.
+2021; Lin et al.
+2019; Wang and He 2021).
+As a standard defense method, adversarial training (Zhang et al.
+2020, 2019; Grathwohl et al.
+2020; Wang et al.
+2023b) aims to enhance the robustness of the target model by training it with adversarial examples.
+Another branch of adversarial defense involves purifying the data stream (Salman et al.
+2020; Hill, Mitchell, and Zhu 2020; Yoon, Hwang, and Lee 2021; Nie et al.
+2022) to remove potential adversarial perturbations or noise that could deceive the model.
+In fact, all of the above methods have shown limitations in robustness, resulting in a lower classification performance on both adversarial and clean data.
+In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
+Continual Learning
+Continual learning (CL) (Liu et al.
+2022; Kirkpatrick et al.
+2017; Wang et al.
+2022; Zhou et al.
+2022) aims to learn from a sequence of new classes without forgetting old ones and attracts a lot of attention to various computer vision tasks.
+In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (called no exemplar) (Zhu et al.
+2022), and some methods have attempted to learn new classes in a few-shot scenario (Zhang et al.
+2021) in which only a small number of new class data is gainable.
+Attempting to break the one-shot defense assumption, a concurrent work (Zhou and Hua 2024) is proposed to alleviate catastrophic forgetting in a simple continual defense scenario.
+In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
+Threat Model and Defense Principles
+The defense system is deployed online and protects the target model against a dynamic environment marked by a plethora of unique attacks.
+As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs.
+Due to the reality, we propose a practical threat model in which distinct attacks emerge across various stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e.
+adversarial examples and their ground-truths).
+Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget.
+The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model.
+The continual defense scenario is formulated as follows.
 Figure 2: Illustration of CAD.
 Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting.
-In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning.
+In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate the overfitting brought by future fine-tuning.
 Then, a weight estimator model $f_{p}$ is trained for model ensemble.
-In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback.
+In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense budget.
 To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process.
-After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high accuracy in classifying clean and adversarial images.
+After all, the classifier is refined, and the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high performance in classifying clean and adversarial data.
 Our scenario is established on the $N$-way $K$-shot classification paradigm.
-Attacks are launched under the gray-box setting [32] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space.
+Attacks are launched under the gray-box setting (Taran et al.
+2019) in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space.
 The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage.
-At the $i$-th stage, where $i=1,2,\cdots,T$, a new attack $A_{i}(\cdot)$.
-is detected and the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack.
+At the $i$-th stage, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$.
+are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack.
 Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
 Few-shot adaptation.
-The increase in defense feedback directly correlates with the frequency of successful attacks on the target model.
-As defenders, we must proactively adapt the defense based on few-shot feedback to prevent potential security disasters that could arise from delayed action following an abundance of feedback.
+The increase in defense budget directly correlates with the frequency of successful attacks on the target model.
+The defense system must proactively adapt the defense based on few-shot budget to prevent potential security disasters that could arise from delayed action.
 Principle 3
 Memory-efficient adaptation.
-Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to memory constraints.
+Over time, the ongoing influx of attacks results in accumulating defense budgets, potentially leading to memory constraints.
 In practical terms, the defender may not have sufficient memory space to accommodate this accumulation.
-Consequently, storing all received defense feedback for adaptation is impractical and should be avoided.
+Consequently, storing all received defense budget for adaptation is impractical and should be avoided.
 Principle 4
-High accuracy in classifying both clean and adversarial images.
-The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation.
-This decline could potentially hurt crucial real-world business operations.
-As a result, a defender must prioritize maintaining performance on both clean and adversarial images.
-4 Continual Adversarial Defense Framework
-As shown in  Fig.
-2, Continual Adversarial Defense (CAD) consists of continual adaptation to defend the target model while preventing catastrophic forgetting ( Sec.
-4.1), embedding reservation to alleviate overfitting brought by the few-shot feedback ( Sec.
-4.2), prototype augmentation for memory efficiency ( Sec.
-4.3), and model ensemble to ensure the classification performance on both clean and adversarial images(Section 4.4).
-4.1 Continual Adaptation to New Attacks
+High accuracy in classifying both clean and adversarial data.
+The previous defense strategies have led to a sacrifice in the classification accuracy of clean data, resulting in performance degradation.
+This decline could potentially harm crucial real-world business operations.
+As a result, a defender must prioritize maintaining performance on both clean and adversarial data.
+Continual Adversarial Defense Framework
+As shown in Figure 2, Continual Adversarial Defense (CAD) framework consists of continual adaptation to defend the target model while preventing catastrophic forgetting, embedding reservation to alleviate overfitting brought by the few-shot budget, prototype augmentation for memory efficiency, and model ensemble to ensure the classification performance on both clean and adversarial data.
+Continual Adaptation to New Attacks
 In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting.
-In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$.
+In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$.
 The defense model is designed to tackle adversarial examples, complementing the target model.
-In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$.
-Inspired by continual learning (CL) [27], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
-Following [22], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage.
-The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion.
+In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to the new attack using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$.
+Inspired by continual learning (CL) (Rebuffi et al.
+2017), we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
+Due to the internal similarity of some adversarial attacks (Wang et al.
+2023a), trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion.
+Old attacks occurring in new stages also lead to unnecessary adaptation.
+Therefore, we conduct model tests and filter out the useless data from the defense budget that can be correctly classified.
+Following (Liu et al.
+2022), we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage.
+The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$.
 Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^1] We omit the networks’ parameter in formulas:
 $y^{i}=y+N\times{i}.$ (1)
-The initial defense model $f_{0}$ will be trained in Sec.
-4.2.
-After this, we freeze the feature extractor $\varphi$ and fine-tune the part of the classifier expanded for the new attack $A_{i}$ using few-shot feedback $\mathcal{A}_{\mathrm{train}}^{i}$ in Sec.
-4.3.
+The initial defense model $f_{0}$ will be trained in the Embedding Reservation Section.
+After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in the Prototype Augmentation Section.
 For evaluation, we are just focusing on which class the image is placed in rather than being threatened by which attack. Thus the prediction for an instance $\mathbf{x}$ is:
 $y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\%N$ (2)
-4.2 Embedding Reservation for Few-Shot Adaptation
-A problem brought by the few-shot feedback in Principle 2 is over-fitting which is common in few-shot learning [30].
+Embedding Reservation for Few-Shot Adaptation
+A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning (Snell, Swersky, and Zemel 2017).
 To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
-First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as "virtual classes" [43], where $V=N\times T$ is the number of virtual classes, i.e.
+First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ”virtual classes” (Zhang et al.
+2021), where $V=N\times T$ is the number of virtual classes, i.e.
 the reserved classes for future attacks.
 Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
-Second, virtual instances are constructed by manifold mixup [35]:
+Second, virtual instances are constructed by manifold mixup (Verma et al. 2019):
 $\mathbf{z}=h_{2}(\lambda h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\lambda)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
-where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as [47].
+where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as (Zhou et al.
+2022).
 $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e.
 $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
-Third, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
-$\mathcal{L}_{\mathrm{v}}=$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{ce}(f_{0}(\mathbf{z}),\hat{y})+$
-$+$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}}).$ (4)
+Finally, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
+$\mathcal{L}_{0}=$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})+$
+$+$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}})+F_{\mathrm{ce}}(f_{0}(\mathbf{z}),\hat{y}).$ (4)
 where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label.
 $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes.
-$\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss [46], and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
+$\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss (Zhang and Sabuncu 2018), and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
 $\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})=f_{0}(\mathbf{x}_{\mathrm{adv}}^{0})\otimes(\mathbf{1}-\mathrm{onehot}(y^{0}))$ (5)
-where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector. The final loss for training the initial defense model is:
-$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\quad\mathcal{L}_{0}=\mathcal{L}_{\mathrm{ce}}+\gamma\mathcal{L}_{\mathrm{virtual}}$ (6)
-In Sec.
-4.2, it forces an instance to be away from the reserved virtual classes (item 1) and pushes the mixed instance towards a virtual class (item 2) and away from other classes (item 3).
-Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved [47].
-The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot feedback.
-4.3 Prototype Augmentation for Memory Efficiency
+where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector.
+In Eq.
+Embedding Reservation for Few-Shot Adaptation, it pushes an instance away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces a virtual instance away from known class (item 3) while towards virtual class (item 4).
+Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved (Zhou et al.
+2022).
+The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
+Prototype Augmentation for Memory Efficiency
 In response to Principle 3, we use prototype augmentation to efficiently manage memory.
-When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50].
-Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [27].
-However, preserving old feedback may also lead to memory shortage.
-For memory efficiency, we adopt prototype augmentation [51] to maintain the decision boundary of previous stages, without saving any feedback.
+When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased (Zhu et al.
+2021).
+Many CL methods store a fraction of old data to jointly train the model with current data to address this issue (Rebuffi et al.
+2017).
+However, preserving old budget may also lead to memory shortage.
+For memory efficiency, we adopt prototype augmentation (Zhu et al.
+2022) to maintain the decision boundary of previous stages, without saving any budget.
 We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
 $\mathbf{p}_{B}=U_{B}(P_{e}),\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (6)
-where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ the same as ProtoNet [30]:
+where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ following ProtoNet (Snell, Swersky, and Zemel 2017):
 $\mathbf{p}_{e}^{j}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (7)
-where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of feedback for each class.
-In each stage, the defense model is adapted using the following loss:
+where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of budget for each class.
+At each stage, the defense model is adapted using the following loss:
 $\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{proto}}$ (8)
-4.4 Model Ensemble for Clean and Adversarial Image Classification
-To maintain high classification accuracy on both clean and adversarial images mentioned in Principle 4, we propose the model ensemble as the last part of CAD.
-Ensemble adversarial training (EAT) [34] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean images.
-We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models
-We adopt the self-perturbation [37] to train the weight estimator at the first stage.
+Model Ensemble for Clean Data Classification
+To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD.
+Ensemble adversarial training (EAT) (Tramèr et al.
+2018) which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data.
+We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
+We adopt self-perturbation (Wang et al.
+2023a) to train the weight estimator at the first stage.
 Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial.
-Details of the self-perturbation are given in the supplementary.
 The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
 $\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (9)
-In this way, both the clean images and the adversarial images are correctly classified.
-The overall algorithm of CAD framework is presented in Algorithm 1.
+In this way, both clean and adversarial images are correctly classified.
+For the overall algorithm of the CAD framework, details about self-perturbation, and limitations of CAD, please refer to the supplementary materials.
+Experiment
+To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams.
+For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) against each attack after adaptation for defense performance evaluation.
+(2) Average accuracy of all attacks occurring within and before each stage, i.e.
+average incremental accuracy (AIAcc) (Rebuffi et al.
+2017), for assessing the knowledge retention of the model.
+Datasets
+Experiments are conducted on CIFAR-10 and ImageNet-100.
+CIFAR-10 (Krizhevsky, Hinton et al.
+2009) contains 5000 images for training and 1000 images for testing in each of the 10 classes.
+ImageNet-100 (Tian, Krishnan, and Isola 2020) is a subset of ImageNet (Deng et al.
+2009) and contains 100 classes with 1000 images for training and 100 images for testing in each class.
+Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively.
+Before training, the above datasets endured data augmentation including horizontal flip and random crop.
+Attack Algorithms
+We choose PGD-$l_{\infty}$ (Madry et al.
+2017) as the initial attack.
+After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM (Kurakin, Goodfellow, and Bengio 2016), RFGSM (Tramèr et al.
+2018), MIM (Dong et al.
+2018), DIM (Wu et al.
+2021), NIM (Lin et al.
+2019), SNIM (Lin et al.
+2019), VNIM and VMIM (Wang and He 2021) to compose the attack pool in defense performance evaluation.
+The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al.
+2019).
+As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM and SNIM, $l_{2}$ attack CW (Carlini and Wagner 2017) and DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016), $l_{1}$ attack EAD (Chen et al.
+2018) and EADEN (Chen et al.
+2018), and $l_{0}$ attack OnePixel (Su, Vargas, and Sakurai 2019) and SparseFool (Modas, Moosavi-Dezfooli, and Frossard 2019) compose the attack pool.
+Details of attack parameters are given in the supplementary.
+Stage:Attack 0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM Clean
+None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
+TRADES (Zhang et al. 2019) 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
+GAIRAT (Zhang et al. 2020) 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
+RPF (Dong and Xu 2023) 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
+DMAT (Wang et al. 2023b) 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
+ADP (Yoon, Hwang, and Lee 2021) 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
+DiffPure (Nie et al. 2022) 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
+CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
+Table 1: Classification accuracy(%) against attacks($l_{\infty}$) on ImageNet-100.
+Volumn ”Clean” represents the standard accuracy of clean images.
+$\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD in case no constraints on memory and defense budget, and is considered to be the empirical upper bound marked with a gray background.
+Stage:Attack 0:PGD 2:SNIM 3:VNIM 4:CW 5:DeepF. 6:EAD 7:EADEN 8:OnePixel 9:SparseF.
+None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
+EWC (Kirkpatrick et al. 2017) 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
+SSRE (Zhu et al. 2022) 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
+BEEF (Wang et al. 2022) 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
+FACT (Zhou et al. 2022) 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
+CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
+Table 2: Continual defense performance (average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10.
+DeepF.
+and SparseF.
+denote DeepFool and SparseFool respectively.
+Implementation Details
+Both the target model and defense model use ResNet (He et al.
+2016) as the backbone network in our method.
+Specifically, we use WideResNet-28-10 for CIFAR-10 following (Grathwohl et al.
+2020) and ResNet-50 for ImageNet-100 following (Laidlaw, Singla, and Feizi 2020).
+A 4-layer ConvNet is used as the weight estimator model.
+The cosine classifier (Gidaris and Komodakis 2018) is adopted as the classifier of the defense model.
+We utilize Torchattacks (Kim 2020) to generate adversarial images.
+Shots of defense budget per class is set to $K=10$.
+The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack).
+Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following (Zhou et al.
+2022).
+Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
+Defense Baselines
+We first compare with adversarial training methods TRADES (Zhang et al.
+2019), GAIRAT (Zhang et al.
+2020), RPF (Dong and Xu 2023) and DMAT (Wang et al.
+2023b), and purification methods ADP (Yoon, Hwang, and Lee 2021) and DiffPure (Nie et al.
+2022) in defense performance evaluation.
+As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC (Kirkpatrick et al.
+2017), SSRE (Zhu et al.
+2022), BEEF (Wang et al.
+2022), and FACT (Zhou et al.
+2022).
+Besides, we propose another variant of our method as the empirical upper bound, denoted as premium-CAD.
+For premium-CAD, there is no constraint of memory and no concerns about the amount of defense budget.
+We train additional attack-specific defense models using abundant defense budget in each stage, and ensemble them with the target model by the weight estimator.
+For details about the implementation of defense baseline and premium-CAD, please refer to supplementary materials.
+Comparisons to Baselines
+To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks.
+For CAD, the defense environment is dynamic, with attacks occurring at different stages.
+The other baseline methods operate in a static defense environment that does not differentiate the order of attacks.
+As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack.
+When compared with the top baseline DMAT and ADP on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack.
+This underscores its capacity for continuous adaptation in response to emerging attacks.
+Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$.
+This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
+(a) PGD
+(b) CW
+(c) EAD
+(d) OnePixel
+Figure 3: T-SNE diagrams of features of test data encoded by the defense model after all stages.
+Each color represents a class of CIFAR10.
+After several steps of adaptation, CAD retains the capability to categorize the attacked data.
+To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods on CIFAR-10 in a dynamic scenario formed by attacks under different norms $l_{p=0,1,2,\infty}$.
+As depicted in Table 2, CAD maintains robust defense performance despite changes in attack types.
+However, as the variety of attack types increases, the defense performance gradually diminishes.
+This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense.
+Compared with the top baseline method FACT, CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
+It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense budget and memory resources, premium-CAD emerges as the superior choice.
+Ablation Study
+As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting.
+Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
+Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel Clean
+w/o $\mathcal{L}_{0}$ 94.5 82.4 76.1 71.8 95.8
+w/o $\mathcal{L}_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
+w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
+CAD(ours) 94.5 89.9 86.6 85.3 95.8
+Table 3: AIAcc(%) on CIFAR-10 for ablation.
+Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and OnePixel($l_{0}$) to validate the efficacy of each component.
+As shown in Table 3, CAD without $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage.
+The absence of model ensemble leads to the inability to uphold performance on clean images.
+Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Figure 3 to visualize the defense capability of CAD after adaptation.
+These features are encoded by the defense model after the last stage of adapting to OnePixel attack.
+From (a) to (d) in Figure 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types.
+After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
+Multiple Attacks in One Stage
+Stage 0 1 2 3
+$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
+$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
+$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
+Table 4: Acc(%) against multiple attacks in one stage.
+In practice, it is common for multiple attacks to occur simultaneously within a single stage.
+To assess CAD’s adaptability in such scenarios, on CIFAR-10, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks per stage.
+As presented in Table 4, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively.
+The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
+Number of Shots for Adaptation
+Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel
+$K=15$ 94.5 90.0 88.4 85.5
+$K=10$ 94.5 89.9 86.6 85.3
+$K=5$ 94.5 89.4 86.1 84.6
+$K=1$ 94.5 87.9 84.6 83.4
+Table 5: AIAcc(%) with different budget scale.
+As indicated in Principle 2, the defense adaptation relies on few-shot defense budget.
+To investigate the influence of the budget shots $K$, we conducted tests with $K=15,10,5,1$ on CIFAR-10.
+The results presented in Table 5 illustrate that defense performance diminishes as $K$ decreases.
+Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($83.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
+Memory Allowance for Adaptation
+As outlined in Principle 3, we need to employ memory efficiently for continual defense.
+In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
+Cache Size Model Size
+Stage 1 2 3 1 2 3
+CAD 50.3K 75.5K 100.4K 141M 141M 141M
+$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
+Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
+stage on CIFAR-10.
+As depicted in Table 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size.
+This underscores our approach’s ability to utilize efficient memory.
+Conclusion
+The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks.
+In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework.
+CAD is engineered to collect few-shot defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages.
+In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot budget adaptation, memory-efficient adaptation, and high accuracy for both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles.
+Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
+Appendix
 Algorithm 1 Continual Adversarial Defense
-Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
+Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot budget $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
 Output: Defense model ensemble.
@@
 8:     Generate prototypes of $A_{i}$ using Eq.7 and add them to prototype set $P_{e}$.
-9:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq. 15.
-10:  end for
-5 Experiment
-To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from three research streams.
-For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy against each attack after adaptation for defense performance evaluation.
-(2) Average accuracy of all attacks occurring within and before each stage, i.e.
-average incremental accuracy [27], for assessing the knowledge retention of the model.
-Datasets
-CIFAR-10 [17] is a widely used dataset for adversarial attack and defense, containing 5000 images for training and 1000 images for testing in each of the 10 classes.
-ImageNet-100 [33] is a subset of ImageNet [4] and contains 100 classes with 1000 images for training and 100 images for testing in each class.
-Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively.
-Before training, the above datasets endured data augmentation including horizontal flip and random crop.
-Attack Algorithms
-We choose PGD-$l_{\infty}$ [23] as the initial attack.
-After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM [18], RFGSM [34], MIM [6], DIM [40], NIM [21], SNIM [21], VNIM [38] and VMIM [38] to compose the attack pool in defense performance evaluation.
-The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44].
-As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM [38] and SNIM [21], $l_{2}$ attack CW [1] and DeepFool [25], $l_{1}$ attack EAD [2] and EADEN [2], and $l_{0}$ attack OnePixel [31] and SparseFool [24] compose the attack pool.
-Details of attack parameters are given in the supplementary.
-Stage 0 1 2 3 4 5 6 7 8 -
-Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
-None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
-TRADES [44] 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
-JEM [10] 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
-PAT [19] 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
-GAIRAT [45] 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
-FastAdv [14] 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
-RPF [5] 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
-DMAT [39] 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
-EBM [12] 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
-ADP [41] 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
-DiffPure [26] 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
-CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
-$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
-Table 1: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10.
-Volumn ”Clean” represents the standard accuracy of clean images.
-$\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense feedback.
-Implementation Details
-Both the target model and defense model use ResNet as the backbone network in our method.
-Specifically, we use WideResNet-28-10 [42] for CIFAR-10 following [10] and ResNet-50 [11] for ImageNet-100 following [19].
-A 4-layer ConvNet is used as the weight estimator model.
-The cosine classifier [8] is adopted as the classifier of the defense model.
-We utilize Torchattacks [15] for generating adversarial images.
-The amount of feedback per class is set to $K=10$.
-The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack).
-Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [47].
-Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
-Defense Baselines
-We first compare to adversarial training methods TRADES [44], JEM [10], GAIRAT [45], FastAdv [14], RPF [5] and DMAT [39], and purification methods EBM [12], ADP [41] and DiffPure [26] in defense performance evaluation.
-As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC [16], SSRE [51], BEEF [36], and FACT [47].
-Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback.
-In premium-CAD, we train additional attack-specific defense models using abundant feedback in each stage, and ensemble them with the target model by the weight estimator.
-Details about the implementation of defense baseline and premium-CAD are given in the supplementary.
-5.1 Comparisons to Baselines
-To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks.
-For CAD, the defense environment is dynamic, with attacks occurring at different stages.
-The other baseline methods operate in a static defense environment that does not differentiate the order of attacks.
-All the evaluation is conducted on the whole test set of CIFAR-10 and ImageNet-100.
-Classification accuracy is reported in Tab.
-1 and Tab.
-2.
-Stage 0 1 2 3 4 5 6 7 8 -
-Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
-None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
-TRADES [44] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
-GAIRAT [45] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
-RPF [5] 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
-DMAT [39] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
-ADP [41] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
-DiffPure [26] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
-CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
-$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
-Table 2: Classification accuracy(%) against various attacks($l_{\infty}$) on ImageNet-100.
-Stage 0 1 2 3 4 5 6 7 8
-Attack PGD SNIM VNIM CW DeepF. EAD EADEN OnePixel SparseF.
-None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
-EWC [16] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
-SSRE [51] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
-BEEF [36] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
-FACT [47] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
-CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
-$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
-Table 3: Continual defense performance (Average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10.
-DeepF.
-and SparseF.
-denote DeepFool [25] and SparseFool [24] respectively.
-As depicted in Tab.
-1 and Tab.
-2, despite the small scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack.
-When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ on CIFAR-10 and $6\%$ on ImageNet-100 for each attack.
-This underscores its capacity for continuous adaptation in response to emerging attacks.
-Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy on CIFAR-10 and $88.1\%$ on ImageNet-100, while the target model stands at $96.4\%$ on CIFAR-10 and $89.8\%$ on ImageNet-100.
-This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
-To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods through experiments on CIFAR-10 with attacks under different norms $l_{p=0,1,2,\infty}$.
-All baseline methods are functioned within our dynamic scenario.
-As depicted in Tab.
-3, both CAD and premium-CAD maintain robust defense performance despite changes in attack types.
-However, as the variety of attack types increases, the defense performance gradually diminishes.
-This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense.
-Compared with the top baseline method FACT [47], CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
-It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense feedback and memory resources, premium-CAD emerges as the superior choice.
-(a) PGD
-(b) CW
-(c) EAD
-(d) OnePixel
-Figure 3: T-SNE diagrams of features extracted from adversarial test data encoded by the defense model after all stages.
-Each color represents a class of CIFAR10.
-After several steps of adaptation, CAD retains the capability to categorize the attacked data.
-5.2 Ablation Study
-As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting.
-Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
-Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and
-Stage 0 1 2 3 -
-Attack PGD CW EAD OnePixel Clean
-w/o $L_{\mathrm{virtual}}$ 94.5 82.4 76.1 71.8 95.8
-w/o $L_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
-w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
-CAD(ours) 94.5 89.9 86.6 85.3 95.8
-Table 4: Average incremental accuracy(%) on CIFAR-10 for ablation.
-OnePixel($l_{0}$) to validate the efficacy of each component.
-As shown in Tab.
-4, CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage.
-The absence of model ensemble leads to the inability to uphold performance on clean images.
-Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Fig.
-3 to visualize the defense capability of CAD after adaptation.
-These features are encoded by the defense model after the last stage of adapting to OnePixel attack.
-From (a) to (d) in Fig.
-3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types.
-After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
-5.3 Multiple Attacks in One Stage
-In practice, it is common for multiple attacks to occur simultaneously within a
-Stage 0 1 2 3
-$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
-$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
-$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
-Table 5: Classification accuracy(%) about the number of attacks($l_{\infty}$) in one stage.
-single stage.
-To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks on CIFAR-10.
-As presented in Tab.
-5, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively.
-The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
-5.4 Number of Shots for Adaptation
-As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we
-Stage 0 1 2 3
-Attack PGD CW EAD OnePixel
-$K=15$ 94.5 90.0 88.4 85.5
-$K=10$ 94.5 89.9 86.6 85.3
-$K=5$ 94.5 89.4 86.1 84.6
-$K=1$ 94.5 87.9 84.6 83.4
-Table 6: Average incremental accuracy(%) with different feedback scale.
-conducted tests with $K=15,10,5,1$ on CIFAR-10.
-The results presented in Tab.
-6 illustrate that defense performance diminishes as $K$ decreases.
-Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($93.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
-5.5 Memory Allowance for Adaptation
-As outlined in Principle 3, we need to employ memory efficiently for continual defense.
-In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
-Cache Size Model Size
-Stage 1 2 3 1 2 3
-CAD 50.3K 75.5K 100.4K 141M 141M 141M
-$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
-Table 7: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
-stage on CIFAR-10.
-As depicted in Tab.
-7, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size.
-This underscores our approach’s ability to utilize efficient memory.
-6 Conclusion
-The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks.
-In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework.
-CAD is engineered to collect few-shot defense feedback from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages.
-In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles.
-Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
-A Weight Estimator Base on Self-Perturbation
-In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation [37] to train a weight estimator to ensemble the target model and the defense model.
+9:     Refine the classifier using cached prototypes $P_{e}$.
+10:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq.9.
+11:  end for
+Appendix A More Details about CAD
+The overall algorithm of the CAD framework is presented in Algorithm 1.
+Model Ensemble Base on Self-Perturbation
+In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation (Wang et al.
+2023a) to train a weight estimator to ensemble the target model and the defense model.
 We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples.
@@
 15:  return $\mathbf{x}^{p}$
-Self-Perturbation
 As presented in Algorithm 2, we perturb each point in the point-wise pattern and each block in the block-wise pattern in a stochastic direction, where blocks are random neighborhoods of a set of scattered points.
@@
 $\mathbf{x}^{p}=\mathbf{x}^{r}+\boldsymbol{\eta}^{p}.$ (11)
-Model Ensemble
 Trained with self-perturbation, the weight estimator $f_{p}$ is able to distinguish between the clean and adversarial images.
@@
 In this way, both the clean images and the adversarial images are correctly classified.
-B Premium-CAD
+Stage 0 1 2 3 4 5 6 7 8 -
+Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
+None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
+TRADES (Zhang et al. 2019) 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
+JEM (Grathwohl et al. 2020) 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
+PAT (Laidlaw, Singla, and Feizi 2020) 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
+GAIRAT (Zhang et al. 2020) 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
+FastAdv (Jiang et al. 2023) 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
+RPF (Dong and Xu 2023) 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
+DMAT (Wang et al. 2023b) 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
+EBM (Hill, Mitchell, and Zhu 2020) 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
+ADP (Yoon, Hwang, and Lee 2021) 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
+DiffPure (Nie et al. 2022) 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
+CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
+$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
+Table 7: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10.
+Volumn ”Clean” represents the standard accuracy of clean images.
+$\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense budget.
+Appendix B Premium-CAD
 We propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback.
@@
 3:     Train a stage-evolved weight estimator $f_{p}^{i}$.
-4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Eq. 15.
+4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Equation 14.
 5:  end for
+Stage 0 1 2 3 4 5 6 7 8
+Attack VMIM VNIM NIM DIM MIM RFGSM BIM SNIM PGD
+CIFAR 94.8 94.6 93.4 93.5 94.2 93.4 92.5 92.2 93.3
+ImageNet 82.5 82.1 82.0 82.0 92.4 82.2 81.9 82.3 81.7
+Table 8: Classification Acc (%) against changed attack sequence.
 Stage-Evolved Weight Estimator
@@
 The overall algorithm of premium-CAD framework is presented in Algorithm 3. The loss function for training attack-specific defense models and stage-evolved weight estimators is Cross-Entropy Loss.
-C Implementation Details
-C.1 Adversarial Attack
-In this work, all of the adversarial attacks are implemented using TorchAttacks [15].
-The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44].
+Appendix C More Implementation Details
+Adversarial Attack
+In this work, all adversarial attacks are implemented using TorchAttacks (Kim 2020).
+The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al.
+2019).
 Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively.
-The number of perturbed pixels of OnePixel [31] is set to $50$.
+The number of perturbed pixels of OnePixel (Su, Vargas, and Sakurai 2019) is set to $50$.
 The other parameters are the default ones in TorchAttacks.
-C.2 Defense Baselines
-The adversarial training baselines are implemented using RobustBench [3], while purification baselines are implemented using open-source code and model checkpoints.
-CL baselines are implemented using PyCIL [48].
+Defense Baselines
+The adversarial training baselines are implemented using RobustBench (Croce et al.
+2021), while purification baselines are implemented using open-source code and model checkpoints.
+CL baselines are implemented using PyCIL (Zhou et al.
+2023).
 The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100.
-CL models are trained under the same settings as CAD.
+CL models are trained in the same settings as in CAD.
 Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.
+Appendix D More Experimental Results
+Continual Adversarial Defense on CIFAR-10
+We compare our CAD with more adversarial defense baselines on CIFAR-10.
+As presented in Table 7, both CAD and premium-CAD demonstrate strong adaptability to each type of attack.
+When compared with the top baseline methods DMAT (Wang et al.
+2023b) and ADP (Yoon, Hwang, and Lee 2021), CAD outperforms them by approximately $10\%$ for each attack.
+As for clean images, CAD achieves $95.8\%$ accuracy while the target model stands at $96.4\%$.
+Change the sequential of attacks
+For different attack sequences, we reverse the sequential of attacks and conduct an experiment.
+As shown in Table 8, when the attack sequence changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack order.
+Change the intensity of attacks
+Stage 0 1 2 3 4 5 6
+ImageNet-$\epsilon/255$ $4$ $6$ $8$ $10$ $12$ $14$ $16$
+Acc 83.7 83.6 83.3 83.3 83.4 82.2 82.5
+CIFAR-$\epsilon/255$ $8$ $10$ $12$ $14$ $16$ $18$ $20$
+Acc 94.5 94.4 94.4 93.9 94.0 93.2 93.4
+Table 9: Classification Acc (%) against various attack intensity of PGD($l_{\infty}$).
+For different attack intensities, we conduct experiments using PGD($l_{\infty}$) with various $\epsilon$.
+As shown in Table 9, when the attack intensity changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack intensity.
+Averaged accuracy when the number of classes increases
+We present curves depicting the relationship between averaged accuracy and the number of classes in Figure 4.
+As the number of classes increases, CAD continues to perform well on both the new attack and previous attacks.
+Figure 4: Average Acc with 10 classes and 100 classes per stage for CIFAR10 and ImageNet100.
+Appendix E Limitations and Future Work
+The increase in attacks denotes the expansion of the defense model classifier, especially when tackling large datasets such as the original ImageNet.
+Although CAD refines the defense budget before each stage’s classifier expansion, too many classes can lead to an undesirable decrease in classification performance.
+In the future, we will focus on this issue and refine the classifier to guarantee the consistent performance of the defense model.